OpenStudy (anonymous):

(-5,1) and -4x+3y=-6

OpenStudy (anonymous):

what are you trying to find

Nnesha (nnesha):

do you have to find parallel line perpendicular line or do you want to know if this point ( -5 ,1) is a solution ???

OpenStudy (emmamink):

Because it is not a solution. Welcome to Open Study! I am Emma. I am here to help guide you to the answers you need, so you don't come back 5min later and ask a similar question. You can also tag me. Just put @EmmaMink

OpenStudy (anonymous):

slope intercept form

OpenStudy (emmamink):

Oh, ok!

OpenStudy (emmamink):

Slope intercept form looks like this: y = mx + b, where m is the slope and b is the y-intercept point!

OpenStudy (emmamink):

So if we have -4x + 3y = -6, we need to put it in that form. -6 is the b and -4 is m. Ok, so we can add 4x to each side to isolate the y, 3y = -4x - 6. Now, we can divide both sides by three....and we have y = -4/3x - 2. Now plug in those coordinates you have and see if it works. It doesn't.

OpenStudy (emmamink):

medal if u want!

OpenStudy (texaschic101):

slope intercept form : y = mx + b -4x + 3y = -6 ---- add 4x to both sides 3y = 4x - 6 -- divide by 3 y = 4/3x - 2 That is your equation in slope intercept form.